# Direct imaging of hydrogen-driven dislocation and strain field evolution in a stainless steel grain

### v1

@@ 2.4 Dislocation strain field evolution
Figure 5: Evolution of the internal strain field surrounding the large dislocation. Three time points are presented as rows in a,b,c,d). a), Translucent morphologies of the grain with dislocation networks. Slices through $\varepsilon_{111}$ are in-plane to the Burgers vector and the normal to the $\left(\bar{1}11\right)$ plane. The average lattice parameter is listed for each time. b), A $\varepsilon_{111}$ slice capturing a section of the large dislocation with pure edge character. See Supporting Video S3 for $\varepsilon_{111}$ slices through the entire grain. c), Theoretical model devoid of hydrogen, $\varepsilon_{111,\mathrm{model}}$, of b) (Methods). d), Circular line profiles, drawn at a 30 nm radius from each dislocation core, compared to the model. The shaded region is the standard deviation of the experimental values in the neighboring 26 pixels. Line profiles start at the Burgers vector (horizontal) and run anticlockwise. e), Circular line profiles drawn for each time point of the experiment, plotted as a surface indicating $\varepsilon_{111}$ values. f), Top-down view of e). g), Maximum and minimum values of the line profiles, averaged over a range of $\pi/4$. The shaded region corresponds to one standard deviation of each value.
@@
The evolution of strain along the circular line profiles during charging is shown in Figure 5e and f. Figure 5g shows how the maximum and minimum values, averaged over ±$\pm\pi/8$ around each maximum/minimum, evolve during charging. We observe a reduction of the elastic strain surrounding the dislocation by up to 35% between the start and end of hydrogen charging. This relative reduction agrees well with the computed reduction of edge dislocation-carbon atom interaction energy in $\alpha$-Fe upon introduction of hydrogen [46]. Our observations reveal hydrogen elastic shielding (Figure 5) which has been predicted by simulations [16], and indirectly inferred from in situ TEM observations of dislocation pileup evolution following the introduction of hydrogen [27]. These direct BCDI measurements of subtle changes in dislocation strain field are important, since they provide fundamental validation of hydrogen-induced dislocation stress shielding, which is core to the HELP mechanism.
@@ 3 Conclusion
Using 316 SS as a model system, our results capture the evolution of a single dislocation undergoing glide relaxation, followed by climb. We propose that this sequence is initiated by the introduction of hydrogen into the lattice, which greatly enhances dislocation mobility via several complementary mechanisms: A hydrogen-induced reduction of lattice friction, or Peierls stress, has been predicted by density functional theory (DFT) studies on pure metals using the Peierls-Nabarro model [47]. However, in 316 SS, this reduction is small compared to the pinning stress from alloying elements, given that 316 SS’s yield stress is more than twice that of pure iron. In addition, hydrogen accumulation has been predicted to lower the dislocation core energy [48] and reduce the dislocation strain field [46], as we have directly measured (Figure 5). This reduces the elastic interactions of the dislocation with the surrounding environment, such as other dislocations and point defects, allowing easier dislocation glide. Furthermore, our observation that hydrogen greatly enhances dislocation climb mobility (Figure 4) suggests that also during glide, dislocations will be able to “climb around” obstacles that would otherwise lead to dislocation pinning. Together, these observations suggest that, in the presence of hydrogen, the currently accepted view of glide-dominated dislocation motion [16, 14, 27] may no longer hold.
Our findings showcase the unique capability of in situ BCDI to monitor the evolution of nanoscale deformation and defects in a grain exposed to hydrogen under bulk conditions. This advancement is particularly timely as many third-generation synchrotrons are being upgraded to fourth-generation facilities, offering greater coherent flux for BCDI experiments with enhanced spatial and temporal resolution. Importantly, in situ BCDI is applicable to most alloy systems and perfectly complements higher-resolution electron microscopy techniques, computational models, and atomistic simulations to shed light on HE degradation pathways. Understanding these internal mechanisms is key to engineering alloys with enhanced resistance to hydrogen-induced degradation, which are urgently needed to enable the green hydrogen economy.
@@ 4.1 Sample preparation
The SS sample was produced using high pressure torsion (HPT) to produce sub-micron grain sizes appropriate for BCDI. A sheet of 316 A4 austenitic stainless steel (composition shown in Table LABEL:supp-table:composition), 1 mm thick, was obtained from RS Components (stock number 264-7241) and cut into a 5 mm diameter disk. The disk then went through HPT to refine the grain microstructure. HPT was performed in a quasi-constrained set-up [36] on a Zwick Roell Z100 Materials Testing machine. Using 5 mm diameter anvils, a compressive force of 80 kN was applied to the face of the disk and maintained for 30 turns using a rotational speed of 5 °$\mathrm{s^{-1}}$. The final disk thickness after HPT was 300 $\mu$m. The disk was then annealed in vacuum ($5\times 10^{-6}$ mbar) at 700 °C for 1 hour. Following annealing, the disk was sequentially ground with 800, 1200, and 2500 grit SiC paper, followed by polishing in 3 $\mu$m and 1 $\mu$m diamond suspension.
@@ 4.3 Hydrogen charging
A 1 L hydrogen charging solution was prepared using 4 g of NaOH (0.1 mol/L) and 5 g thiourea (0.07 mol/L) dissolved in deionized water, yielding a pH of 13. After an appropriate grain was found, the hydrogen charging solution was continuously pumped into the cell at 5 mL/min. Repeated $111$ CXDPs from the grain were measured in solution, without charging, over three hours to verify that there was no effect on the sample. Next, hydrogen charging was performed using chronopotentiometry with a two electrode setup, keeping the current fixed at 0.2 mA (0.5 mA/cm2). During the charging, two repeated BCDI measurements for the $111$ reflection were measured approximately every 40 minutes, with each scan requiring approximately 10 minutes to complete. We noticed bubbles slowly form during charging.
@@ 4.5 BCDI strain calculations
where $a_{0,\mathrm{ref.}}$ is the reference lattice parameter, determined at -3.4 h.
@@ 4.10 Dislocation strain field modeling
The $\varepsilon_{111,\mathrm{model}}$ for the large dislocation was created using the dislocation node positions determined from BCDI, the Burgers vector (using the local lattice parameter, $a_{0,t}$), and $\nu=0.28$. Each dislocation node was joined by an edge, which was then connected to another point to form a dislocation triangle. The displacement field for this triangular dislocation loop was determined using the solution developed by Barnett [68, 69], which was then numerically differentiated to determine the lattice strain field, $\varepsilon_{\mathrm{model}}$,
@@ Acknowledgements
F.H. thanks Thomas D. Swinburne and Max Boleininger for insightful comments. D.Y., G.H., and F.H. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 714697). K.S. acknowledges funding from the General Sir John Monash Foundation. A.M. acknowledges funding from EDF (Électricité de France). Work at Brookhaven National Laboratory was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Contract No. DESC0012704. Work performed at UCL was supported by EPSRC. The authors acknowledge the use of the Advanced Research Computing (ARC) facility at the University of Oxford [70]. Experiments were performed at the Advanced Photon Source, a US Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357.
@@
Conceptualisation: D.Y., F.H.; Data Curation: D.Y., M.S.; Formal Analysis: D.Y., M.S., E.T., I.K.R., F.H.; Funding Acquisition: E.T., I.K.R., F.H.; Investigation: D.Y., G.H., K.S., R.J.H, W.C.; Methodology: D.Y., B.D., K.S., R.J.H, W.C.; Project Administration: F.H.; Resources: D.Y., G.H., K.S., A.M, F.H.; Software: D.Y., M.S., E.T., F.H.; Supervision: E.T., I.K.R., F.H.; Validation: D.Y., M.S.; Visualisation: D.Y., M.S.; Writing – Original Draft Preparation: D.Y., M.S., F.H.; Writing – Review & Editing: All authors.

### v2

@@ 2.4 Dislocation strain field evolution
Figure 5: Evolution of the internal strain field surrounding the large dislocation. Three time points are presented as rows in a,b,c,d). a), Translucent morphologies of the grain with dislocations. Slices through $\varepsilon_{111}$ are in-plane to the Burgers vector and the normal to the $\left(\bar{1}11\right)$ plane. The average lattice parameter is listed for each time. b), A $\varepsilon_{111}$ slice capturing a section of the large dislocation with pure edge character. See Supporting Video S3 for $\varepsilon_{111}$ slices through the entire grain. c), Theoretical elastic model, $\varepsilon_{111,\mathrm{model}}$, of the large dislocation devoid of hydrogen (Methods). d), Circular line profiles, drawn at a 30 nm radius from each dislocation core, compared to the model. The shaded region is the standard deviation of the experimental values in the neighboring 26 pixels. Line profiles start at the Burgers vector (horizontal) and run anticlockwise. e), Circular line profiles drawn for each time point of the experiment, plotted as a surface indicating $\varepsilon_{111}$ values. f), Top-down view of e). g), Maximum and minimum values of the line profiles, averaged over a range of $\pi/4$. The shaded region corresponds to one standard deviation of each value. The model values are offset to overlap with the experimental values at negative times.
@@
The evolution of strain along the circular line profiles during charging is shown in Figure 5e and f. Figure 5g shows how the maximum and minimum values, averaged over ±$\pm\pi/4$ around each maximum/minimum, evolve during charging. We observe a reduction of the elastic strain surrounding the dislocation by up to 35% between the start and end of hydrogen charging. This relative reduction agrees well with the computed reduction of edge dislocation-carbon atom interaction energy in $\alpha$-Fe upon the introduction of hydrogen [59].
At the nanoscale, we attribute this elastic strain reduction to the formation of a hydrogen atmosphere in the vicinity of the dislocation. Dislocations are relatively large and strong hydrogen trapping sites ($\sim 27\ \mathrm{kJ\ mol^{-1}}$) [22], so the hydrogen atoms surrounding the dislocations are relatively immobile at room temperature [24]. Therefore, the evolution of this localized accumulation can be probed using the heterogeneous strain. Figure 5g also shows that the maximum $\varepsilon_{111}$ deviates more from the elastic hydrogen-free model than the minimum $\varepsilon_{111}$ This suggests that hydrogen prefers to reside in the tensile region of the dislocation [16]. Alternatively, the reduction in tensile strain could also be due to the hydrogen-facilitated diffusion of solutes, resulting in solute segregation [53], but this is unlikely since we are at room temperature.
Our observations reveal hydrogen elastic shielding (Figure 5) which has been predicted by simulations [16], and indirectly inferred from in situ TEM observations of dislocation pileup evolution following the introduction of hydrogen [41]. These direct BCDI measurements of subtle changes in dislocation strain field are important, since they provide fundamental validation of hydrogen-induced dislocation stress shielding, which is core to the HELP mechanism.
2.5 Proposed mechanisms
Our results capture the evolution of a single dislocation undergoing glide relaxation, followed by climb. We propose that this sequence is initiated by the introduction of hydrogen into the lattice, which greatly enhances dislocation mobility via several complementary mechanisms. A hydrogen-induced reduction of lattice friction, or Peierls stress, has been predicted by DFT studies on pure metals using the Peierls-Nabarro model [60]. However, in 316 SS, this reduction is small compared to the pinning stress from alloying elements, given that 316 SS’s yield stress is more than twice that of pure iron. In addition, hydrogen accumulation has been predicted to lower the dislocation core energy [61] and reduce the dislocation strain field [59], as we have directly measured (Figure 5). This reduces the elastic interactions of the dislocation with the surrounding environment, such as other dislocations and point defects, allowing easier dislocation glide. Furthermore, our observation that hydrogen greatly enhances dislocation climb mobility (Figure 4) suggests that also during glide, dislocations will be able to “climb around” obstacles that would otherwise lead to dislocation pinning. Together, these observations suggest that, in the presence of hydrogen, the currently accepted view of glide-dominated dislocation motion [16, 14, 41] may no longer hold.
@@ 3 Conclusion
Our findings showcase the unique capability of in situ BCDI to monitor the evolution of nanoscale deformation and defects in a grain exposed to hydrogen under bulk conditions. This advancement is particularly timely as third-generation synchrotrons are being upgraded to fourth-generation facilities, offering orders of magnitude greater coherent flux for BCDI experiments that will enhance spatial and temporal resolution. This could enable exciting possibilities for exploring hydrogen induced evolution in adjacent grains or extended grains using Bragg ptychography [62] In situ BCDI is applicable to most alloy systems and perfectly complements higher-resolution electron microscopy techniques, computational models, and atomistic simulations to shed light on HE degradation pathways. Understanding these internal mechanisms is key to designing alloys with enhanced resistance to hydrogen-induced degradation, such as engineering micro- and nano-structures to mitigate dislocation movement or vacancy formation. These hydrogen-resistant materials are urgently needed to enable the green hydrogen economy.
@@ 4.1 Sample preparation
The SS sample was produced using high pressure torsion (HPT) to produce sub-micron grain sizes appropriate for BCDI. A sheet of 316 A4 austenitic stainless steel (composition shown in Table LABEL:supp-table:composition), 1 mm thick, was obtained from RS Components (stock number 264-7241) and cut into a 5 mm diameter disk. The disk then went through HPT to refine the grain microstructure. HPT was performed in a quasi-constrained set-up [50] on a Zwick Roell Z100 Materials Testing machine. Using 5 mm diameter anvils, a compressive force of 80 kN was applied to the face of the disk and maintained for 30 turns using a rotational speed of 5 ° $\mathrm{s^{-1}}$. The final disk thickness after HPT was 300 $\mu$m. The disk was then annealed in vacuum ($5\times 10^{-6}$ mbar) at 700 °C for 1 hour, ramping up and down at 4 °C $\mathrm{min}^{-1}$. Following annealing, the disk was sequentially ground with 800, 1200, and 2500 grit SiC paper, followed by polishing in 3 $\mu$m and 1 $\mu$m diamond suspension.
Powder X-ray diffraction (Figure LABEL:supp-fig:powder_XRD) shows that after HPT, the sample will be predominately composed of deformation-induced martensite. The heat treatment removes macroscale residual stresses and the sample undergoes recovery and recrystallization to form austenite [63, 64]. The measured austenite grain was likely formed during heat treatment due to the relatively low misorientation and very low observed strain (relative to the average lattice parameter of the grain) near the surface of the grain. We note that carbides could have formed during the annealing, which could attract solutes that have become mobile due to hydrogen.
@@ 4.3 Hydrogen charging
A 1 L hydrogen charging solution was prepared using 4 g of NaOH (0.1 mol $\mathrm{L}^{-1}$) and 5 g thiourea (0.07 mol $\mathrm{L}^{-1}$) dissolved in deionized water, yielding a pH of 13. After an appropriate grain was found, the hydrogen charging solution was continuously pumped into the cell at 5 mL $\mathrm{min}^{-1}$. Repeated $111$ CXDPs from the grain were measured in solution, without hydrogen charging, over three hours to verify that there were no X-ray effects on the sample. Furthermore, any X-ray induced changes to the strain field are negligible (Supporting Text). Next, hydrogen charging was performed using chronopotentiometry with a two electrode setup, keeping the current fixed at 0.2 mA (0.5 mA $\mathrm{cm^{-2}}$). During the charging, two repeated BCDI measurements for the $111$ reflection were measured approximately every 40 minutes, with each scan requiring approximately 10 minutes to complete. We noticed bubbles slowly form during charging.
@@ 4.5 BCDI strain calculations
where $a_{0,\mathrm{ref.}}$ is the reference lattice parameter, determined at -3.4 h. $a_{0}$ was calculated using Bragg’s law and the $2\theta$ angle determined from the angular position of the diffractometer.
@@ 4.10 Dislocation strain field modeling
We assume that there are no externally applied stresses. The $\varepsilon_{111,\mathrm{model}}$ for the large dislocation was created using the dislocation node positions determined from BCDI, the Burgers vector (using the local lattice parameter, $a_{0,t}$), and $\nu=0.28$. Each dislocation node was joined by an edge, which was then connected to another point to form a dislocation triangle. The displacement field for this triangular dislocation loop was determined using the solution developed by Barnett [84, 85], which was then numerically differentiated to determine the lattice strain field, $\varepsilon_{\mathrm{model}}$,
@@ Acknowledgements
F.H. thanks Thomas D. Swinburne and Max Boleininger for insightful comments. D.Y., G.H., and F.H. acknowledge funding from the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 714697). K.S. acknowledges funding from the General Sir John Monash Foundation. A.M. acknowledges funding from EDF (Électricité de France). Work at Brookhaven National Laboratory was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Contract No. DESC0012704. This research used the Materials Synthesis and Characterization Facility of the Center for Functional Nanomaterials (CFN), which is a U.S. Department of Energy Office of Science User Facility, at Brookhaven National Laboratory under Contract No. DE-SC0012704. Experiments were performed at the Advanced Photon Source, a US Department of Energy (DOE) Office of Science User Facility operated for the DOE Office of Science by Argonne National Laboratory under Contract No. DE-AC02-06CH11357. Work performed at UCL was supported by EPSRC. The authors acknowledge the use of the Advanced Research Computing (ARC) facility at the University of Oxford [86].
@@
Conceptualization: D.Y., F.H.; Data Curation: D.Y., M.S.; Formal Analysis: D.Y., M.S., E.T., I.K.R., F.H.; Funding Acquisition: E.T., I.K.R., F.H.; Investigation: D.Y., G.H., K.S., R.J.H., W.C., D.N.; Methodology: D.Y., B.D., K.S., R.J.H., W.C.; Project Administration: F.H.; Resources: D.Y., G.H., K.S., A.M, F.H.; Software: D.Y., M.S., E.T., F.H.; Supervision: E.T., I.K.R., F.H.; Validation: D.Y., M.S.; Visualisation: D.Y., M.S.; Writing – Original Draft Preparation: D.Y., M.S., F.H.; Writing – Review & Editing: All authors.